# Point Cloud Matters: Rethinking the Impact of Different Observation Spaces on Robot Learning

### v2

@@ 4.2 Study on performance of different observations with and without pre-training (Q1, Q2)
| PickCube | 0.17 | 0.24 | 0.01 | 0.00 | 0.00 | 0.71 | 0.70 | 0.04 | 0.01 | |
| StackCube | 0.03 | 0.00 | 0.00 | 0.00 | 0.00 | 0.04 | 0.00 | 0.00 | 0.00 | |
| TurnFaucet | 0.08 | 0.07 | 0.00 | 0.00 | 0.00 | 0.32 | 0.36 | 0.28 | 0.00 | |
| Peg- Insertion- Side | Grasp | 0.78 | 0.45 | 0.06 | 0.80 | 0.46 | 0.82 | 0.83 | 0.06 | 0.02 |
| Align | 0.07 | 0.02 | 0.00 | 0.01 | 0.02 | 0.09 | 0.16 | 0.00 | 0.00 | |
| Insert | 0.01 | 0.00 | 0.00 | 0.00 | 0.00 | 0.01 | 0.01 | 0.00 | 0.00 | |
| Excavate | 0.01 | 0.02 | 0.00 | 0.00 | 0.00 | 0.17 | 0.24 | 0.02 | 0.00 | |
| Hang | 0.52 | 0.42 | 0.00 | 0.00 | 0.00 | 0.67 | 0.72 | 0.65 | 0.09 | |
| Pour | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| Fill | 0.36 | 0.04 | 0.03 | 0.00 | 0.01 | 0.21 | 0.68 | 0.21 | 0.02 | |
@@
| Mean S.R. | 0.13 | 0.08 | 0.03 | 0.05 | 0.06 | 0.20 | 0.23 | - | - | |
| Mean Rank | 2.68 | 3.63 | 3.84 | 4.26 | 3.42 | 2.00 | 1.58 | - | - | |
@@ 4.3.2 Zero-Shot Generalization to Visual Changes
We further discuss the problem of visual generalization. The StackCube task in ManiSkill2 was selected for its complexity and apparent preference for the RGB modality, as demonstrated by its higher initial accuracy. We consider 3 kinds of visual changes. 1) Lighting: We vary lighting intensity, testing levels at , from the default intensity of . 2) Noise: We introduce visual noise by switching the rendering mode from rasterization to ray tracing, disabling the denoiser, and varying the number of ray tracing samples per pixel to , and . 3) Background color: We alter the original gray floor to red and green, varying the ’R’ or ’G’ values to respectively. Illustrations of these visual changes are displayed in Fig. 3, Fig. 3 and Fig. 3. The results are plotted in Fig. 2. For comprehensive results on all our generalization experiments, please refer to Tab. 7, Tab. 8, Tab. 9, Tab. 10 and Tab. 11 in Appendix D.
@@ 4.4 Study on sample efficiency (Q4)
To evaluate the sample efficiency across different observation spaces, we conducted experiments on RLBench tasks with reduced training data. Specifically, each method was trained using only and out of the total training trajectories. The results are presented in Tab. 5 and Appendix E.
@@ 4.5 Study on design decisions on point cloud observation space (Q5)
In this section, we demonstrate that not only the point cloud itself matters but also how it is utilized is equally, if not more, important. Through extensive experimentation on ManiSkill2 tasks, we aim to provide insights into the use of point clouds in robotic learning problems. First, we investigate the importance of coordinate and color information. Next, we examine the influence of point cloud sampling strategies, particularly the widely adopted FPS sampling, used to obtain a fixed number of points for convenient input to policies, as seen in [67, 77]. Additionally, we explore the use of pointmap, a novel format that stacks RGB images with explicit coordinate information. The results of these experiments are detailed in Tab. 6. Full results of each task can be found in Appendix F.
@@ Appendix C Implementation Details
General Implementations. Our implementation leverages PyTorch [54], a powerful framework for deep learning applications. For convenient and flexible experimental configuration, we build our codebase upon PyTorch Lightning [19] and Hydra [70]. For the ResNet50 encoder, we utilized the official model available in TorchVision [48], aligning with the implementation in R3M. This approach ensures consistency with the original work regarding weights and feature extraction. We meticulously adhered to the original implementations for the ViT and VC-1 models. In our experiments, we chose the ViT-base and corresponding VC-1-base models to maintain parameter consistency with other methods used in our study, promoting relative fairness. In the case of MultiViT and MultiMAE, our implementation is based on the official MultiViT-B model. We employed the strongest pre-trained weights in the official repository, pre-trained simultaneously on RGB, depth, and semantic modalities, enhancing the model’s robustness and versatility. For SpUNet and PonderV2, we adopt the SpUNet34 architecture as described in the official PonderV2 paper, ensuring alignment with their proposed methodology and optimizing the model’s performance in handling point cloud data.
@@
Appendix D Full Zero-Shot Generalization Results
@@
D.1 Zero-Shot Generalization to Camera View
@@ D.2 Zero-Shot Generalization to Visual Changes
The results of visual changes are listed in Tab. 11.
Appendix E Full Sample Efficiency Results
Full sample efficiency results on RLBench are listed in Tab. 12.
@@ Appendix E Full Sample Efficiency Results
Appendix F Full Design Decision Results
Full design decision results on ManiSkill2 are listed in Tab. 13.

### v3

@@ 4.2 Study on performance of different observations with and without pre-training (Q1, Q2)
| PickCube | 0.17 | 0.24 | 0.34 | 0.58 | 0.00 | 0.71 | 0.70 | 0.04 | 0.01 | |
| StackCube | 0.03 | 0.00 | 0.59 | 0.03 | 0.00 | 0.04 | 0.00 | 0.00 | 0.00 | |
| TurnFaucet | 0.08 | 0.07 | 0.24 | 0.30 | 0.00 | 0.32 | 0.36 | 0.28 | 0.00 | |
| Peg- Insertion- Side | Grasp | 0.78 | 0.45 | 0.94 | 0.68 | 0.46 | 0.82 | 0.83 | 0.06 | 0.02 |
| Align | 0.07 | 0.02 | 0.11 | 0.03 | 0.02 | 0.09 | 0.16 | 0.00 | 0.00 | |
| Insert | 0.01 | 0.00 | 0.01 | 0.00 | 0.00 | 0.01 | 0.01 | 0.00 | 0.00 | |
| Excavate | 0.01 | 0.02 | 0.23 | 0.03 | 0.00 | 0.17 | 0.24 | 0.02 | 0.00 | |
| Hang | 0.52 | 0.42 | 0.77 | 0.56 | 0.00 | 0.67 | 0.72 | 0.65 | 0.09 | |
| Pour | 0.00 | 0.00 | 0.06 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| Fill | 0.36 | 0.04 | 0.72 | 0.03 | 0.01 | 0.21 | 0.68 | 0.21 | 0.02 | |
@@
| Mean S.R. | 0.13 | 0.08 | 0.23 | 0.13 | 0.06 | 0.20 | 0.23 | - | - | |
| Mean Rank | 3.37 | 4.47 | 2.37 | 4.00 | 4.21 | 2.42 | 1.89 | - | - | |
@@ 4.3.2 Zero-Shot Generalization to Visual Changes
We further discuss the problem of visual generalization. The StackCube task in ManiSkill2 was selected for its complexity and apparent preference for the RGB modality, as demonstrated by its higher initial accuracy. We consider 3 kinds of visual changes. 1) Lighting: We vary lighting intensity, testing levels at , from the default intensity of . 2) Noise: We introduce visual noise by switching the rendering mode from rasterization to ray tracing, disabling the denoiser, and varying the number of ray tracing samples per pixel to , and . 3) Background color: We alter the original gray floor to red and green, varying the ’R’ or ’G’ values to respectively. Illustrations of these visual changes are displayed in Fig. 3, Fig. 3 and Fig. 3. The results are plotted in Fig. 2. For comprehensive results on all our generalization experiments, please refer to Tab. 12, Tab. 13, Tab. 14, Tab. 15 and Tab. 16 in Appendix E.
@@ 4.4 Study on sample efficiency (Q4)
To evaluate the sample efficiency across different observation spaces, we conducted experiments on RLBench tasks with reduced training data. Specifically, each method was trained using only and out of the total training trajectories. The results are presented in Tab. 5 and Appendix F.
@@ 4.5 Study on design decisions on point cloud observation space (Q5)
In this section, we demonstrate that not only the point cloud itself matters but also how it is utilized is equally, if not more, important. Through extensive experimentation on ManiSkill2 tasks, we aim to provide insights into the use of point clouds in robotic learning problems. First, we investigate the importance of coordinate and color information. Next, we examine the influence of point cloud sampling strategies, particularly the widely adopted FPS sampling, used to obtain a fixed number of points for convenient input to policies, as seen in [73, 83]. Additionally, we explore the use of pointmap, a novel format that stacks RGB images with explicit coordinate information. The results of these experiments are detailed in Tab. 6. Full results of each task can be found in Appendix G.
@@
5 Additional Ablations
In this section, we further provide additional ablation experiments.
Ablation study on encoder architecture. We assigned XYZ coordinates to each image pixel based on its UV coordinates, treating images as "plain point clouds" for processing by point cloud networks. We experimented with both RGB and RGB-D modalities using PointNet and SpUNet, applying both ACT and diffusion policies. The results, shown in Tables 8 and 8, strongly support our claims.
| PointNet | SpUNet | ||||||
| Tasks | RGB | RGB-D | PCD | RGB | RGB-D | PCD | |
| PickCube | 0.680 | 0.195 | 0.843 | 0.352 | 0.282 | 0.740 | |
| StackCube | 0.252 | 0.093 | 0.348 | 0.068 | 0.052 | 0.220 | |
| TurnFaucet | 0.393 | 0.335 | 0.000 | 0.447 | 0.268 | 0.388 | |
| Peg- Insertion- Side | Grasp | 0.675 | 0.565 | 0.765 | 0.658 | 0.668 | 0.810 |
| Align | 0.308 | 0.095 | 0.395 | 0.220 | 0.132 | 0.280 | |
| Insert | 0.027 | 0.000 | 0.005 | 0.007 | 0.013 | 0.008 | |
| Excavate | 0.035 | 0.038 | 0.268 | 0.058 | 0.097 | 0.032 | |
| Hang | 0.820 | 0.853 | 0.828 | 0.728 | 0.817 | 0.840 | |
| Pour | 0.112 | 0.110 | 0.135 | 0.090 | 0.000 | 0.095 | |
| Fill | 0.615 | 0.147 | 0.905 | 0.842 | 0.827 | 0.660 | |
| Mean S.R. | 0.392 | 0.243 | 0.449 | 0.347 | 0.316 | 0.407 | |
| PointNet | SpUNet | ||||||
| Tasks | RGB | RGB-D | PCD | RGB | RGB-D | PCD | |
| PickCube | 0.863 | 0.430 | 0.900 | 0.623 | 0.567 | 0.710 | |
| StackCube | 0.495 | 0.140 | 0.238 | 0.143 | 0.120 | 0.035 | |
| TurnFaucet | 0.133 | 0.298 | 0.365 | 0.308 | 0.308 | 0.318 | |
| Peg- Insertion- Side | Grasp | 0.855 | 0.550 | 0.868 | 0.805 | 0.807 | 0.815 |
| Align | 0.243 | 0.072 | 0.285 | 0.143 | 0.072 | 0.093 | |
| Insert | 0.013 | 0.000 | 0.008 | 0.003 | 0.005 | 0.000 | |
| Excavate | 0.195 | 0.235 | 0.278 | 0.060 | 0.100 | 0.168 | |
| Hang | 0.740 | 0.712 | 0.778 | 0.540 | 0.600 | 0.673 | |
| Pour | 0.095 | 0.040 | 0.125 | 0.038 | 0.005 | 0.000 | |
| Fill | 0.252 | 0.165 | 0.693 | 0.203 | 0.195 | 0.208 | |
| Mean S.R. | 0.388 | 0.264 | 0.454 | 0.286 | 0.278 | 0.302 | |
Ablation study on diffusion policy variants. By default, we chose the UNet version of the diffusion policy for its stability and robustness. Many subsequent works, like [61, 83, 73], also use this variant. To validate this more clearly, we have conducted additional validations to compare UNet and transformer versions of diffusion policy under different observation spaces on the PickCube task. As shown in Tab. 10, the point cloud performs the best under different policy variants.
Ablation study on point cloud coordinate frame. In Frame Mining [49], the authors discovered that using the end-effector (EE) frame can lead to even better results. We conduct additional experiments with the EE frame using both PointNet and SpUNet. The results are in the Tab. 10. We found that the EE frame can indeed enhance point cloud performance in many cases, though this can vary depending on specific tasks and networks.
| ACT Policy | Diffusion Policy | ||||||||
| PointNet | SpUNet | PointNet | SpUNet | ||||||
| Tasks | World | EE | World | EE | World | EE | World | EE | |
| PickCube | 0.843 | 0.915 | 0.740 | 0.540 | 0.900 | 0.935 | 0.740 | 0.842 | |
| StackCube | 0.348 | 0.442 | 0.220 | 0.465 | 0.238 | 0.145 | 0.220 | 0.370 | |
| TurnFaucet | 0.000 | 0.000 | 0.388 | 0.000 | 0.365 | 0.545 | 0.388 | 0.435 | |
| Peg- Insertion- Side | Grasp | 0.765 | 0.647 | 0.810 | 0.873 | 0.868 | 0.890 | 0.810 | 0.910 |
| Align | 0.395 | 0.195 | 0.280 | 0.245 | 0.285 | 0.140 | 0.280 | 0.175 | |
| Insert | 0.005 | 0.007 | 0.008 | 0.015 | 0.008 | 0.001 | 0.008 | 0.000 | |
| Excavate | 0.268 | 0.005 | 0.032 | 0.025 | 0.278 | 0.278 | 0.113 | 0.245 | |
| Hang | 0.828 | 0.810 | 0.840 | 0.825 | 0.778 | 0.820 | 0.798 | 0.770 | |
| Pour | 0.135 | 0.172 | 0.095 | 0.096 | 0.125 | 0.140 | 0.095 | 0.000 | |
| Fill | 0.905 | 0.923 | 0.660 | 0.558 | 0.693 | 0.740 | 0.660 | 0.068 | |
| Obs. Space | Encoder | Variant | Mean S.R. |
| RGB | ResNet50 | UNet | 0.165 |
| RGB-D | ResNet50 | UNet | 0.340 |
| Point Cloud | PointNet | UNet | 0.900 |
| Point Cloud | SpUNet | UNet | 0.740 |
| RGB | ResNet50 | Transformer | 0.280 |
| RGB-D | ResNet50 | Transformer | 0.340 |
| Point Cloud | PointNet | Transformer | 0.915 |
| Point Cloud | SpUNet | Transformer | 0.527 |
6 Real-World Experiments
As a benchmark work, our goal is to ensure our project is reproducible, open-source, easy to follow, and accessible to all researchers. In addition, some recent literature such as [46] indicates that the experimental conclusions in modern simulation environments are consistent with those in the real world. Thus we believe simulators can play a crucial role in providing fair and comparable evaluations. Besides, we also admit the importance of real-world validation. To address this, we conducted additional real-world experiments using the open-sourced low-cost-robot [43] equipped with two Intel RealSense D415 RGB-D cameras. Our bimanual setups (including 2 leader arms, 2 follower arms, 2 cameras, etc.) cost about $2000 in total, making them affordable and easy to replicate for researchers.
| Obs. Space | Reach Cube | Pick Cube | Fold Cloth |
| RGB | 0.60 | 0.05 | 0.65 |
| RGB-D | 0.30 | 0.20 | 0.50 |
| Point Cloud | 0.80 | 0.40 | 0.80 |
We’ve also open-sourced our real-world codebase [13] for easy reproduction by other researchers. We designed three tasks: 1) Reach Cube: A single arm with rigid objects. The robot is required to reach towards a cube and touch it. 2) Pick Cube: A single arm with rigid objects. The robot is required to pick up the cube and hold it in the air. 3) Fold Cloth: Two dual arms with soft-body objects. The robot is required to simultaneously catch one side of the cloth with two grippers and then fold the cloth in half. The detailed setups, our workstation, and task visualizations are shown in Appendix D. The real-world results shown in Tab. 11 align with our simulated experiments, further supporting our conclusions.
@@ 8 Conclusion and Limitations
Acknowledgments and Disclosure of Funding
This work is funded in part by the National Key R&D Program of China No.2022ZD0160102, and Shanghai Artificial Intelligence Laboratory.
@@ Appendix C Implementation Details
General Implementations. Our implementation leverages PyTorch [59], a powerful framework for deep learning applications. For convenient and flexible experimental configuration, we build our codebase upon PyTorch Lightning [21] and Hydra [76]. For the ResNet50 encoder, we utilized the official model available in TorchVision [53], aligning with the implementation in R3M. This approach ensures consistency with the original work regarding weights and feature extraction. We meticulously adhered to the original implementations for the ViT and VC-1 models. In our experiments, we chose the ViT-base and corresponding VC-1-base models to maintain parameter consistency with other methods used in our study, promoting relative fairness. In the case of MultiViT and MultiMAE, our implementation is based on the official MultiViT-B model. We employed the strongest pre-trained weights in the official repository, pre-trained simultaneously on RGB, depth, and semantic modalities, enhancing the model’s robustness and versatility. For SpUNet and PonderV2, we adopt the SpUNet34 architecture as described in the official PonderV2 paper, ensuring alignment with their proposed methodology and optimizing the model’s performance in handling point cloud data. The point cloud-related operations are largely adapted from Pointcept [12].
@@
Appendix D Real-World Setups
In real-world experiments, we implement three modalities using ResNet50 encoders for RGB and RGB-D images, and PointNet for point cloud inputs. The ACT policy is employed across all experiments. All modalities share the same settings and hyper-parameters. For data collection, leader arms are used to teleoperate the follower arms. Two RealSense RGB-D cameras capture the visual observations. Training is conducted on a single NVIDIA A100 GPU with a learning rate of .
Reach Cube: Trained with 45 demonstrations over 100 epochs.
Pick Cube: Trained with 45 demonstrations over 500 epochs.
Fold Cloth: Trained with 50 demonstrations over 1000 epochs.
All models are evaluated with 20 rollouts in the same environment. We report the success rates below, and corresponding videos are available at Google Drive (https://drive.google.com/drive/folders/1UiFgHv9QUPEM2is-N10IJ47DjeYiiFEm?usp=drive_link). The real-world results from these tasks align with our simulated experiments, further supporting our conclusions.
Appendix E Full Zero-Shot Generalization Results
@@
E.1 Zero-Shot Generalization to Camera View
E.2 Zero-Shot Generalization to Visual Changes
The results of visual changes are listed in Tab. 16.
Appendix F Full Sample Efficiency Results
Full sample efficiency results on RLBench are listed in Tab. 17.
Appendix G Full Design Decision Results
Full design decision results on ManiSkill2 are listed in Tab. 18.
@@ Appendix G Full Design Decision Results
| PointNet | SpUNet | ||||||
| Tasks | RGB | RGB-D | PCD | RGB | RGB-D | PCD | |
| PickCube | 0.680 | 0.195 | 0.843 | 0.352 | 0.282 | 0.740 | |
| StackCube | 0.252 | 0.093 | 0.348 | 0.068 | 0.052 | 0.220 | |
| TurnFaucet | 0.393 | 0.335 | 0.000 | 0.447 | 0.268 | 0.388 | |
| Peg- Insertion- Side | Grasp | 0.675 | 0.565 | 0.765 | 0.658 | 0.668 | 0.810 |
| Align | 0.308 | 0.095 | 0.395 | 0.220 | 0.132 | 0.280 | |
| Insert | 0.027 | 0.000 | 0.005 | 0.007 | 0.013 | 0.008 | |
| Excavate | 0.035 | 0.038 | 0.268 | 0.058 | 0.097 | 0.032 | |
| Hang | 0.820 | 0.853 | 0.828 | 0.728 | 0.817 | 0.840 | |
| Pour | 0.112 | 0.110 | 0.135 | 0.090 | 0.000 | 0.095 | |
| Fill | 0.615 | 0.147 | 0.905 | 0.842 | 0.827 | 0.660 | |
| Mean S.R. | 0.392 | 0.243 | 0.449 | 0.347 | 0.316 | 0.407 | |
| PointNet | SpUNet | ||||||
| Tasks | RGB | RGB-D | PCD | RGB | RGB-D | PCD | |
| PickCube | 0.863 | 0.430 | 0.900 | 0.623 | 0.567 | 0.710 | |
| StackCube | 0.495 | 0.140 | 0.238 | 0.143 | 0.120 | 0.035 | |
| TurnFaucet | 0.133 | 0.298 | 0.365 | 0.308 | 0.308 | 0.318 | |
| Peg- Insertion- Side | Grasp | 0.855 | 0.550 | 0.868 | 0.805 | 0.807 | 0.815 |
| Align | 0.243 | 0.072 | 0.285 | 0.143 | 0.072 | 0.093 | |
| Insert | 0.013 | 0.000 | 0.008 | 0.003 | 0.005 | 0.000 | |
| Excavate | 0.195 | 0.235 | 0.278 | 0.060 | 0.100 | 0.168 | |
| Hang | 0.740 | 0.712 | 0.778 | 0.540 | 0.600 | 0.673 | |
| Pour | 0.095 | 0.040 | 0.125 | 0.038 | 0.005 | 0.000 | |
| Fill | 0.252 | 0.165 | 0.693 | 0.203 | 0.195 | 0.208 | |
| Mean S.R. | 0.388 | 0.264 | 0.454 | 0.286 | 0.278 | 0.302 | |
| ACT Policy | Diffusion Policy | ||||||||
| PointNet | SpUNet | PointNet | SpUNet | ||||||
| Tasks | World | EE | World | EE | World | EE | World | EE | |
| PickCube | 0.843 | 0.915 | 0.740 | 0.540 | 0.900 | 0.935 | 0.740 | 0.842 | |
| StackCube | 0.348 | 0.442 | 0.220 | 0.465 | 0.238 | 0.145 | 0.220 | 0.370 | |
| TurnFaucet | 0.000 | 0.000 | 0.388 | 0.000 | 0.365 | 0.545 | 0.388 | 0.435 | |
| Peg- Insertion- Side | Grasp | 0.765 | 0.647 | 0.810 | 0.873 | 0.868 | 0.890 | 0.810 | 0.910 |
| Align | 0.395 | 0.195 | 0.280 | 0.245 | 0.285 | 0.140 | 0.280 | 0.175 | |
| Insert | 0.005 | 0.007 | 0.008 | 0.015 | 0.008 | 0.001 | 0.008 | 0.000 | |
| Excavate | 0.268 | 0.005 | 0.032 | 0.025 | 0.278 | 0.278 | 0.113 | 0.245 | |
| Hang | 0.828 | 0.810 | 0.840 | 0.825 | 0.778 | 0.820 | 0.798 | 0.770 | |
| Pour | 0.135 | 0.172 | 0.095 | 0.096 | 0.125 | 0.140 | 0.095 | 0.000 | |
| Fill | 0.905 | 0.923 | 0.660 | 0.558 | 0.693 | 0.740 | 0.660 | 0.068 | |